# Noise-Tolerant Hybrid Approach for Data-Driven Predictive Control

### v1

@@ 1 Introduction
Indirect methods, exemplified by Subspace Predictive Control (SPC) [8], first identify a prediction model from data, which is subsequently used as an constraint in the optimal control problem. This intermediate modeling phase makes these approaches particularly suitable for integration with large datasets. Such integration allows for reduction in predictor bias without increasing the computational complexity of the design problem [9].
In contrast, direct DPC approaches, such as Data-Enabled Predictive Control (DeePC) [10], instead of estimating a prediction model incorporate the input output samples directly in the optimal control problem [11]. This is possible by introducing a regularization vector "$\mathbf{g}(k)$" based on Willems’ fundamental lemma [12]. This approach enables the predictor to be "control-oriented" specifically focused on generating predictions for control purposes, however, the optimization problem has a larger size as $\mathbf{g}(k)$ is part of the decision variables that needs to be optimized [13].
@@
To improve the noise handling of hybrid approaches, this study proposes a different for estimation of $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ that takes the noise characteristics into consideration. Compared to the mentioned hybrid methods, our approach borrows the assumption of orthogonality of components from [17] and replaces the LQ factorization with SVD which can exctract the system dynamics from the noisy data when estimating the $\mathbf{g}(k)$ components. The idea of using SVD in data driven predictive control is not new as [18, 11, 19] have used this decomposition. In [18], using SVD the length of past data and hence the dimension of Hankel matrix is reduced. [11, 19] propose different applications for SVD in which the regularization vector g can be replaced with a lower dimension vector in direct methods. However, when it comes to hybrid methods, using SVD approach is new and even more important. The fact that factorization is used to select one solution among infinite possible solutions for $\mathbf{g}(k)$ components, makes it crucial to select the solution with more information from the system dynamics rather than the noise. The procedure, importance and efficiency of using SVD in hybrid structure are the highlights of this study.
@@ 2 PRELIMINARIES
Consider a discrete-time linear dynamical system influenced by zero-mean Gaussian noise $\zeta(k)\sim N(0,\sigma^{2}_{\zeta}I)$:
@@
where $x\in\mathbb{R}^{n}$ is the state, $u\in\mathbb{R}^{n_{u}}$ is the control input, $y\in\mathbb{R}^{n_{y}}$ is the measured output, and $A,B,C$ are real matrices of appropriate dimensions. We assume that $(A,B)$ is controllable and $(A,C)$ is observable. By applying a persistently exciting input sequence $\{u(k)\}_{k\in\mathbb{N}[0,T]}$ of length $T$ to system (1), we obtain a corresponding output sequence $\{y(k)\}_{k\in\mathbb{N}[0,T]}$.
For an input-output model, we introduce the past horizon $T_{\text{ini}}$ that limits the window of past input-output data necessary to compute the current output:
$y(k)=\sum_{i=1}^{T_{\text{ini}}}a_{i}y(k-i)+\sum_{i=1}^{T_{\text{ini}}}b_{i}u(k-i),$ (3)
Consider a system whose input and output data can be described, for some real-valued matrix coefficients $a_{i}$ and $b_{i}$. For simplicity of notation, we assume a uniform past data window length, denoted by $T_{\text{ini}}$, for both the input and output signals. Now, for any starting time instant $k\geq 0$ within the available data vector and for any data vector length $j\geq 1$, we define:
@@
$\bar{y}(k,j)$ $:=\operatorname{col}(y(k),\dots,y(k+j-1)).$
Lets define $N$ as the future horizon and using $T_{\text{ini}}$, we construct the following sequences of past and future inputs and outputs:
$\begin{split}&\mathbf{u}_{\text{ini}}(k):=\text{col}(u(k-T_{\text{ini}}),\ldots,u(k-1)),\quad\\ &\mathbf{y}_{\text{ini}}(k):=\text{col}(y(k-T_{\text{ini}}+1),\ldots,y(k)),\\ &\mathbf{u}_{N}(k):=\text{col}(u(k),\ldots,u(k+N)),\\ &\mathbf{y}_{N}(k):=\text{col}(y(k),\ldots,y(k+N)),\end{split}$ (4)
To transition from the model-based prediction to data-driven counterpart, we assume access to an informative set of input-output pairs $\{\bar{u},\bar{y}\}$ of length $T$. Informativity is ensured by the persistent excitation of the input sequence, which must be of order $T_{\text{ini}}+N+n$, implies that $T$ must be sufficiently long, specifically $T\geq n_{u}(T_{\text{ini}}+N+n).$ Let $N\geq T_{\text{ini}}$, the Hankel matrices can be defined as follows:
$\mathbf{U}_{p}$ $,\displaystyle:=[\bar{u}(0,T_{\text{ini}}),$$\dots,\bar{u}(T-1,T_{\text{ini}})],$ (5)
$\mathbf{Y}_{p}$ $,\displaystyle:=[\bar{y}(1,T_{\text{ini}}),$$\dots,\bar{y}(T,T_{\text{ini}})],$
$\mathbf{U}_{f}$ $,\displaystyle:=[\bar{u}(T_{\text{ini}},N),$$\dots,\bar{u}(T_{\text{ini}}+T-1,N)],$
$\mathbf{Y}_{f}$ $,\displaystyle:=[\bar{y}(T_{\text{ini}}+1,N),$$\dots,\bar{y}(T_{\text{ini}}+T,N)].$
which form the foundation of the data-driven predictive control approach.
In the next, we introduce three different DPC strategies, by directly using input–output data instead of an explicit model of the system and a state estimator.
Problem 1 [SPC]: Subspace Predictive Control (SPC), represents an early indirect data-driven MPC technique. It directly incorporates results from subspace identification into the MPC design by estimating prediction matrices from input/output data based on the relationship $\mathbf{y}(k)=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)&\mathbf{y}_{\text{ini}}(k)&\mathbf{u}(k)\end{bmatrix}^{T}.$ The unknown matrix $\mathbf{\Theta}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}$ is estimated by solving the least-squares problem $\min_{\mathbf{\Theta}}\left\|\mathbf{Y}_{f}-\mathbf{\Theta}\begin{bmatrix}\mathbf{U}_{p}&\mathbf{Y}_{p}&\mathbf{U}_{f}\end{bmatrix}^{T}\right\|_{\text{Fro}}^{2}$, with solution $\mathbf{\Theta}^{*}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\hat{\Gamma}}\end{bmatrix}$. Then, the SPC can be formulated as [8]:
@@
Unlike SPC, which relies on estimating prediction matrices, Data-Enabled Predictive Control (DeePC) is a purely data-driven control approach that directly predicts the future output sequence $\mathbf{y}(k)$ from Hankel matrix data. Next, the DeePC problem is defined as follows:
@@
where $\Theta(k)=\{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k),\mathbf{g}(k),\sigma_{y}(k)\}$ and, $l_{g}(\mathbf{g}(k))$ is a user-defined regularization function over $\mathbf{g}(k)$, often chosen as $\|\mathbf{g}(k)\|_{2}^{2}$, leading to the term "DeePC". The variable $\sigma_{y}(k)$ is introduced to prevent infeasibility if $\{\mathbf{u}_{\text{ini}}(k),\mathbf{y}_{\text{ini}}(k)\}$ is not in the behavioral set, especially in the presence of disturbances.
To address the need for regularization in the predictive control problem, a variant of DPC was recently developed [17], employing an optimal (minimum variance) unbiased predictor for future output trajectories.
Problem 3 [SMMPC]: Given immediate past input and measured output signals, find a future input sequence by solving [17]:
@@
The faster execution is attributed to the offline computation of $\mathbf{E}_{1}$, $\mathbf{E}_{2}$, and $\mathbf{E}_{3}$.
@@ 3 Indirect DPC using reduced-order Hankel matrices
Data-driven MPC with Hankel matrices often has more columns than rows, leading to non-unique $\mathbf{g}(k)$, requiring regularization of its norm. DeePC’s large matrices can delay dynamics capture in time-varying systems. Smith et.al, [17] separates $\mathbf{g}(k)$ identification into two phases for a smaller control problem: $\mathbf{g}_{\text{ini}}(k)$ from past data (using LQ factorization for uniqueness due to non-full rank Hankel matrix) and $\mathbf{g}_{\text{f}}(k)$ from future data in the null space of the past Hankel matrix. Inspired by [17], we decompose $\mathbf{g}(k)$ into orthogonal parts:
@@
While sharing the idea of separate $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ identification using past and future data, we address SMMPC’s noise handling limitation. Noise can make Hankel matrices full row rank, causing LQ factorization to prioritize noise over signal, significantly impacting SMMPC performance with noisy data. This study focuses on introducing constraints for solution uniqueness and minimizing sensor noise impact. We propose SVD as a superior factorization method for handling noisy data, detailed in the following section. SMMPC uses LQ factorization to uniquely determine $\mathbf{g}_{\text{ini}}(k)$ from partitioned $\mathbf{U}_{p}$ and $\mathbf{Y}_{p}$, and $\mathbf{g}_{f}(k)$ in their null space using future data. While creative for deterministic cases, this may be suboptimal with noisy data.
We propose a systematic method to handle noise in both the output signals and the Hankel matrices. By leveraging singular value decomposition (SVD), we decompose the Hankel matrices into components associated with system dynamics and noise, respectively, based on the assumption that the noise variance is smaller than the signal variance. This decomposition enables a robust estimation of the system dynamics.
A key contribution of our approach is the replacement of the high-dimensional regularization vector $\mathbf{g}(k)=\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k)$ with lower-dimensional vectors $\eta_{1}(k)$ and $\eta_{2}(k)$, exploiting the rank deficiency of the Hankel matrices. Unlike [17], which uses LQ factorization to separately solve deterministic equations, our method formulates the estimation of $\eta_{1}(k)$ as a constrained optimization problem that simultaneously accounts for deterministic input constraints and stochastic output measurements. This joint optimization is necessary because the deterministic and stochastic components are coupled in our SVD-based formulation. The resulting NTDPC problem is formulated as a convex quadratic program (QP), which is computationally efficient due to offline computation of key matrices and eliminates the need for user-defined regularization parameters, unlike DeePC [20], $\gamma$-DDPC [21, 22], and GDPC [23].
@@ 3.1 Derivation of a Parsimonious Formulation
We condense the mapping in (9) into a minimal parameterization. Consider stacked Hankel matrices representing the past and future input-output sequences:
$\mathbf{Z}_{p}:=\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\end{bmatrix},\quad\mathbf{Z}_{f}:=\begin{bmatrix}\mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix},\quad\mathbf{z}_{\text{ini}}(k):=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\end{bmatrix}$ (10)
Using SVD, we factorize $\mathbf{Z}_{p}$ as:
$\mathbf{Z}_{p}=W\Sigma V^{T}$ (11)
@@
where $W$ and $V$ are unitary matrices. If the system is noise free (i.e. $\zeta=0$), $\Sigma$ is not full rank and we have
$\mathbf{Z}_{p}=\begin{bmatrix}W_{1}&W_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{1}^{T}\\ V_{2}^{T}\end{bmatrix}=L_{1}V_{1}^{T}$ (12)
in which $L_{1}=\begin{bmatrix}W_{1}&W_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}\\ 0\end{bmatrix}$, and $\Sigma_{1}$ is a diagonal matrix containing non-zero singular entries. According to [17], the rank of $\mathbf{Z}_{p}$ is equal to $n_{u}T_{\text{ini}}+n$ and therefore $\Sigma_{1}\in\mathbb{R}^{(n_{u}T_{\text{ini}}+n)\times(n_{u}T_{\text{ini}}+n)}$ and $V_{1}^{T}\in\mathbb{R}^{(n_{u}T_{\text{ini}}+n)\times M_{c}}$, where $M_{c}$ is the number of columns in $\mathbf{Z}_{p}$. (Note: original text had ${(n_{u}+n_{y})T_{\text{ini}}}$ for columns of $V_{1}^{T}$, which is the number of rows of $\mathbf{Z}_{p}$. $V_{1}^{T}$ has $M_{c}$ columns, the same as $\mathbf{Z}_{p}$. Its rows are $n_{u}T_{\text{ini}}+n$. Please verify this dimension.)
Now, from (9) we have:
$\begin{split}\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\Big{(}\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{\text{f}}(k)\Big{)}\end{split}$ (13)
where assuming that $\mathbf{g}_{f}(k)$ is in null space of $\mathbf{Z}_{p}$, one can write
$\begin{split}\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\mathbf{g}_{\text{ini}}(k)\end{split}$ (14)
By substituting $\mathbf{Z}_{p}$ from (12) into (14), yields
$\begin{split}\mathbf{z}_{\text{ini}}(k)=L_{1}V_{1}^{T}\mathbf{g}_{\text{ini}}(k)\end{split}$ (15)
Defining a vector $\eta_{1}(k)\in\mathbb{R}^{n_{u}T_{\text{ini}}+n}$ as $\eta_{1}(k)=V_{1}^{T}\mathbf{g}_{\text{ini}}(k)$, then we will obtain the following least-squares problem as:
$\begin{split}\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k),\end{split}$ (16)
where $\eta_{1}(k)$ will be later calculated by solving an estimation problem. Since $L_{1}$ has full column rank (as $\Sigma_{1}$ is invertible), we can recover $\mathbf{g}_{\text{ini}}(k)$ from $\eta_{1}(k)$ as:
$\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k)$ (17)
In case $\zeta\neq 0$, i.e. $\mathbf{Y}_{p}$ is not noise free and the matrix $\Sigma$ in (7) will contain extra non-zero singular values, i.e.:
$\Sigma=\begin{bmatrix}\Sigma_{1}&0\\ 0&\Sigma_{2}\end{bmatrix}$ (18)
However, assuming that the variance of noise is smaller than the variation of inputs and outputs, the SV decomposition by nature would separate the noise from the signal. Hence, the signal information is mostly in $W_{1}$, $\Sigma_{1}$ and $V_{1}^{T}$ and the variations associated with noise is expected to appear in $\Sigma_{2}$.

### v2

@@ 1 Introduction
Indirect methods, exemplified by Subspace Predictive Control (SPC) [8], first identify a prediction model from data, which is subsequently used as a constraint in the optimal control problem. This intermediate modeling phase makes these approaches particularly suitable for integration with large datasets. Such integration allows for reduction in predictor bias without increasing the computational complexity of the design problem [9].
In contrast, direct DPC approaches, such as Data-Enabled Predictive Control (DeePC) [10], instead of estimating a prediction model incorporate the input output samples directly in the optimal control problem [11]. This is possible by introducing a regularization vector $\mathbf{g}(k)$" based on Willems’ fundamental lemma [12]. This approach enables the predictor to be "control-oriented" specifically focused on generating predictions for control purposes, however, the optimization problem has a larger size as $\mathbf{g}(k)$ is part of the decision variables that needs to be optimized [13].
@@
To improve the noise handling of hybrid approaches, this study proposes a different approach for estimation of $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ that takes into account the noise characteristics. Compared to the mentioned hybrid methods, our approach borrows the assumption of orthogonality of components from [17] and replaces the LQ factorization with SVD which can exctract the system dynamics from the noisy data when estimating the $\mathbf{g}(k)$ components. The application of Singular Value Decomposition (SVD) in data-driven control has recently gained attention, primarily for computational model order reduction within direct control formulations [18, 11, 19]. However, the potential of SVD to geometrically separate signal and noise subspaces within a hybrid predictive framework remains underexplored. This study uses SVD beyond simple dimensionality reduction, utilizing it as a robust filtering mechanism. By embedding this projection into the predictor construction, the proposed NTDPC framework systematically replaces the heuristic regularization typical of DeePC with a rigorous, noise-tolerant estimation step. In [18], using SVD the length of past data and hence the dimension of Hankel matrix is reduced. [11, 19] propose different applications for SVD in which the regularization vector g can be replaced with a lower dimension vector in direct methods. However, when it comes to hybrid methods, using SVD approach is new and even more important. The fact that factorization is used to select one solution among infinite possible solutions for $\mathbf{g}(k)$ components, makes it crucial to select the solution with more information from the system dynamics rather than the noise. The procedure, importance and efficiency of using SVD in hybrid structure are the highlights of this study.
@@ 2 PRELIMINARIES
Consider a discrete-time linear dynamical system influenced by zero-mean Gaussian noise $\zeta(k)\sim\mathcal{N}(0,\Sigma_{\zeta})$:
@@
where $x\in\mathbb{R}^{n}$ is the state, $u\in\mathbb{R}^{n_{u}}$ is the control input, $y\in\mathbb{R}^{n_{y}}$ is the measured output, and $A,B,C$ are real matrices of appropriate dimensions. We assume that $(A,B)$ is controllable and $(A,C)$ is observable. By applying a persistently exciting input sequence $\{u(k)\}_{k=0}^{T-1}$ of length $T$ to system (1), we obtain a corresponding output sequence $\{y(k)\}_{k=0}^{T-1}$. The measurement noise $\zeta(k)$ is assumed to be i.i.d., zero-mean Gaussian, and uncorrelated in time.
For an input-output model, we introduce the past horizon $T_{\text{ini}}$ to characterize the system’s memory. Let $y_{o}(k)$ denote the underlying noise-free output of the system (i.e., the output of (1)-(2) if $\zeta(k)\equiv 0$). The system dynamics and the measurement model are described by:
$y_{o}(k)$ $=\sum_{i=1}^{T_{\text{ini}}}a_{i}y_{o}(k-i)+\sum_{i=1}^{T_{\text{ini}}}b_{i}u(k-i),$ (3a)
$y(k)$ $=y_{o}(k)+\zeta(k).$ (3b)
In this representation, the coefficients $a_{i}$ and $b_{i}$ define the system dynamics. For a general MIMO system, these are real-valued matrices with dimensions $a_{i}\in\mathbb{R}^{n_{y}\times n_{y}}$ and $b_{i}\in\mathbb{R}^{n_{y}\times n_{u}}$, while for a SISO system, they reduce to scalar parameters. Furthermore, for simplicity of notation, we assume the same memory horizon $T_{\text{ini}}$ for both the input and output regressors, rather than assigning distinct lag lengths to $u$ and $y$.
Now, for any starting time instant $k\geq 0$ within the available data vector and for any data vector length $j\geq 1$, we define the trajectory vectors:
@@
$\bar{y}(k,j)$ $:=\operatorname{col}(y(k),\dots,y(k+j-1)).$ (4)
In this notation, the first argument $k$ denotes the starting time instant, and the second argument $j$ explicitly denotes the length of the data window.
Let us define $N$ as the future prediction horizon. Using $T_{\text{ini}}$, we construct the following sequences of past and future inputs and outputs:
$\begin{split}\mathbf{u}_{\text{ini}}(k)&:=\text{col}(u(k-T_{\text{ini}}),\ldots,u(k-1)),\\ \mathbf{y}_{\text{ini}}(k)&:=\text{col}(y(k-T_{\text{ini}}+1),\ldots,y(k)),\\ \mathbf{u}_{N}(k)&:=\text{col}(u(k),\ldots,u(k+N-1)),\\ \mathbf{y}_{N}(k)&:=\text{col}(y(k),\ldots,y(k+N-1)).\end{split}$ (5)
To transition from model-based prediction to a data-driven counterpart, we assume access to an informative set of input-output pairs $\{\bar{u},\bar{y}\}$ of length $T$. Informativity is ensured by the persistent excitation of the input sequence, where the rank of $\mathbf{Z}_{p}$ matrix defined in (11) must be higher than $T_{\text{ini}}+N+n$. This implies that the data length $T$ must be sufficiently long, specifically $T\geq(n_{u}+1)(T_{\text{ini}}+N+n)-1$. The persistency of excitation condition is imposed on the input sequence $\{\bar{u}\}$ , which guarantees that the corresponding Hankel matrix $\mathbf{Z}_{p}$ constructed from the input-output data has sufficient rank.
Let $M=T-T_{ini}-N+1$ be the number of columns in the Hankel matrices. The matrices are defined as:
$\mathbf{U}_{p}$ $=[\bar{u}(0,T_{ini}),\dots,\bar{u}(M-1,T_{ini})],$ (6)
$\mathbf{Y}_{p}$ $=[\bar{y}(0,T_{ini}),\dots,\bar{y}(M-1,T_{ini})],$
$\mathbf{U}_{f}$ $=[\bar{u}(T_{ini},N),\dots,\bar{u}(T_{ini}+M-1,N)],$
$\mathbf{Y}_{f}$ $=[\bar{y}(T_{ini},N),\dots,\bar{y}(T_{ini}+M-1,N)].$
These matrices form the foundation of the data-driven predictive control approach. Next, we introduce three different DPC strategies that utilize this data directly, avoiding the need for an explicit system model or state estimator.
Problem 1 [SPC]: Subspace Predictive Control (SPC) represents an early indirect data-driven MPC technique. It directly incorporates results from subspace identification into the MPC design by estimating prediction matrices from input/output data based on the relationship $\mathbf{y}_{N}(k)=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)^{T}&\mathbf{y}_{\text{ini}}(k)^{T}&\mathbf{u}_{N}(k)^{T}\end{bmatrix}^{T}$. The unknown matrix $\mathbf{\Theta}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}$ is estimated by solving the least-squares problem $\min_{\mathbf{\Theta}}\left\|\mathbf{Y}_{f}-\mathbf{\Theta}\begin{bmatrix}\mathbf{U}_{p}^{T}&\mathbf{Y}_{p}^{T}&\mathbf{U}_{f}^{T}\end{bmatrix}^{T}\right\|_{\text{Fro}}^{2}$, yielding the solution $\mathbf{\Theta}^{*}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\hat{\Gamma}}\end{bmatrix}$. The SPC problem is then formulated as [8]:
@@
Unlike SPC, which relies on an intermediate identification step to find prediction matrices, DeePC is a direct approach that embeds the raw data samples into the optimization problem.
@@
where $\Theta(k)=\{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k),\mathbf{g}(k),\sigma_{y}(k)\}$ and $l_{g}(\mathbf{g}(k))$ is a user-defined regularization function over $\mathbf{g}(k)$, often chosen as $\|\mathbf{g}(k)\|_{2}^{2}$. The variable $\sigma_{y}(k)$ is a slack variable introduced to ensure feasibility in the presence of noise.
While DeePC offers a direct data-driven formulation, the optimization vector $\mathbf{g}(k)$ scales with the dataset size, which can be computationally intensive. To mitigate this and improve noise handling, hybrid approaches have been proposed that pre-compute an optimal predictor offline.
Problem 3 [SMMPC]: A variant of DPC, Signal Matrix Model Predictive Control (SMMPC) [17], employs a minimum-variance unbiased predictor derived offline. The control problem is solved as:
@@
The computational efficiency of SMMPC arises from the offline calculation of the predictor matrices $\mathbf{E}_{1}$, $\mathbf{E}_{2}$, and $\mathbf{E}_{3}$, which replaces the high-dimensional online optimization of $\mathbf{g}(k)$ required in DeePC.
@@ 3 Indirect DPC using reduced-order Hankel matrices
In data-driven predictive control, the Hankel matrices constructed from historical data typically possess significantly more columns than rows ($M\gg(n_{u}+n_{y})T_{ini}$). This structural characteristic renders the linear system for the regularization vector $\mathbf{g}(k)$ underdetermined, admitting an infinite set of possible solutions. Consequently, standard approaches require regularization terms (e.g., minimum norm) to select a specific solution.
To address the underdetermined nature of the problem, we adopt a sequential identification strategy that decomposes the regularization vector $\mathbf{g}(k)$ into two orthogonal components: a past component $\mathbf{g}_{\text{ini}}(k)$ determined by historical data, and a future component $\mathbf{g}_{\text{f}}(k)$ that governs the predictive behavior. This decomposition is formulated as:
@@
Equation (10) poses an identification challenge due to the rank properties of the data matrix. In the noise-free case, the Hankel matrix is rank-deficient (rank equal to the system order plus the input horizon), reflecting the underlying low-dimensional dynamics. However, measurement noise invariably renders the matrix full row rank. In this full-rank scenario, standard approaches like LQ factorization force a unique solution that effectively fits the noise, treating stochastic perturbations as valid system dynamics. To prevent this overfitting, we propose a transformation that enforces the underlying low-rank structure. By assuming a spectral gap, we utilize SVD to separate the dominant signal subspace ($\Sigma_{1}$) from the noise subspace ($\Sigma_{2}$). This allows us to project the high-dimensional $\mathbf{g}(k)$ onto a lower-dimensional latent basis ($\eta_{1},\eta_{2}$), converting the ill-posed noisy problem into a well-posed, noise-tolerant estimation problem.
@@ 3.1 Derivation of a Parsimonious Formulation
We condense the mapping in (10) into a minimal parameterization. Consider the stacked Hankel matrices representing the past and future input-output sequences:
$\mathbf{Z}_{p}:=\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\end{bmatrix},\quad\mathbf{Z}_{f}:=\begin{bmatrix}\mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix},\quad\mathbf{z}_{\text{ini}}(k):=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\end{bmatrix}$. (11)
$Z_{p}$Past Data==
$W$
[…]
$\Sigma_{1}$
$\sigma_{1}$$\sigma_{r}$$\Sigma_{2}\ll$00SignalNoise
$V^{\top}$
[…]
Figure 1: SVD Decomposition of the noisy Hankel matrix $Z_{p}$. $\Sigma_{1}$ captures dominant dynamics; $\Sigma_{2}$ represents noise. $W$ is partitioned into signal columns $W_{1}$ and noise columns $W_{2}$.
By performing SVD factorization on the past data matrix $\mathbf{Z}_{p}$:
$\mathbf{Z}_{p}=W\Sigma V^{T}.$ (12)
@@
where $W$ and $V$ are unitary matrices. In the ideal noise-free case ($\zeta=0$), the matrix $\Sigma$ is rank-deficient. We can partition the decomposition as:
$\mathbf{Z}_{p}=\begin{bmatrix}W_{1}&W_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{1}^{T}\\ V_{2}^{T}\end{bmatrix}=W_{1}\Sigma_{1}V_{1}^{T}=L_{1}V_{1}^{T},$ (13)
where we define $L_{1}:=W_{1}\Sigma_{1}$. Here, $\Sigma_{1}$ is a diagonal matrix containing the non-zero singular values (see Fig.1). The rank of the noiseless $\mathbf{Z}_{p}$ is given by $r=n_{u}T_{\text{ini}}+n$ [21]. Consequently, the dimensions are defined as $\Sigma_{1}\in\mathbb{R}^{r\times r}$ and $V_{1}^{T}\in\mathbb{R}^{r\times M}$, where $M$ is the number of columns in $\mathbf{Z}_{p}$.
In cases of low Signal-to-Noise Ratio (SNR), the singular values of the system dynamics and noise can indeed overlap or "flip," making simple truncation problematic. Therefore, we explicitly clarify that our method operates under the assumption of a sufficient spectral gap between the dominant system dynamics and the noise floor.
Substituting this decomposition into the past trajectory equation (10):
$\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\big(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{\text{f}}(k)\big).$ (14)
We enforce the condition that $\mathbf{g}_{\text{f}}(k)$ lies in the null space of $\mathbf{Z}_{p}$, such that $\mathbf{Z}_{p}\mathbf{g}_{\text{f}}(k)=0$. This simplifies (14) to:
$\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\mathbf{g}_{\text{ini}}(k)=L_{1}V_{1}^{T}\mathbf{g}_{\text{ini}}(k).$ (15)
We now define the latent vector $\eta_{1}(k)\in\mathbb{R}^{r}$ as the projection of the regularization vector onto the active subspace:
$\eta_{1}(k)=V_{1}^{T}\mathbf{g}_{\text{ini}}(k).$ (16)
This yields the following linear system describing the initial trajectory:
$\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k).$ (17)
Since we seek the minimum-norm solution for $\mathbf{g}_{\text{ini}}(k)$ that satisfies the data consistency, we restrict $\mathbf{g}_{\text{ini}}(k)$ to the row space of $\mathbf{Z}_{p}$ (spanned by $V_{1}$). This allows us to recover $\mathbf{g}_{\text{ini}}(k)$ uniquely from $\eta_{1}(k)$ via:
$\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k).$ (18)
In the realistic case where $\zeta\neq 0$, the measured output matrix $\mathbf{Y}_{p}$ is noisy, and $\mathbf{Z}_{p}$ becomes full rank. Let $M$ be the number of columns. By utilizing SVD, yielding a rectangular singular value matrix $\Sigma\in\mathbb{R}^{(n_{u}+n_{y})T_{ini}\times M}$, which is partitioned as:
$\Sigma=\begin{bmatrix}\Sigma_{1}&0\\ 0&\Sigma_{2}\end{bmatrix}$, (19)
where $\Sigma_{1}\in\mathbb{R}^{r\times r}$ contains the $r$ dominant singular values associated with the system dynamics. The second block, $\Sigma_{2}\in\mathbb{R}^{((n_{u}+n_{y})T_{ini}-r)\times(M-r)}$, is a rectangular matrix containing the remaining singular values dominated by noise (padded with zeros as necessary). By truncating $\Sigma_{2}$, we obtain a low-rank approximation that serves as a robust basis for estimating $\eta_{1}(k)$.